OpenStudy (anonymous):

M (3,4) is the midpoint of line JK where J has coordinates(1,2). how do i find the coordinates of K.

OpenStudy (anonymous):

Midpoint refers to the half way point between two points, in your case between point J and K. If J = (1,2) then K = ( x,y) we can work out the mid point formula between them this way Midpoint M = ( [1+x]/2 , [2+y]/2 ) = (3,4) so we have 2 equations a) [1+x]/2 =3 b) [2+y]/2 =4 Solve for x and y we get x=5 y=6 So K = (5,6)